OpenStudy (anonymous):

integral of tan 5Θ dΘ

OpenStudy (anonymous):

I got that 5Θ is your u and the du is 5dΘ but I'm not sure what to do next

OpenStudy (anonymous):

\[\int\tan5\theta~d\theta=\int\frac{\sin5\theta}{\cos5\theta}~d\theta\] Substitute \(u=\cos5\theta\), then \(du=-5\sin5\theta~d\theta\), so the integral becomes \[-\frac{1}{5}\int\frac{du}{u}\]